\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}
\providecommand{\MR}{\relax\ifhmode\unskip\space\fi MR }
% \MRhref is called by the amsart/book/proc definition of \MR.
\providecommand{\MRhref}[2]{%
  \href{http://www.ams.org/mathscinet-getitem?mr=#1}{#2}
}
\providecommand{\href}[2]{#2}
\begin{thebibliography}{10}

\bibitem{chodoshschulzeuniquenessofasymptoticallyconical}
O.~Chodosh and F.~Schulze, \emph{Uniqueness of asymptotically conical tangent
  flows}, Duke Mathematical Journal \textbf{170} (2021), no.~16, 3601--3657.

\bibitem{coldingminicozzidifferentiabilityofthearrivaltime}
T.~H. Colding and W.~P. Minicozzi, \emph{Differentiability of the arrival
  time}, Communications on Pure and Applied Mathematics \textbf{69} (2016),
  no.~12, 2349--2363.

\bibitem{coldingminicozziuniquenesslojasiewicz}
T.~H. Colding and W.~P. Minicozzi~II, \emph{Uniqueness of blowups and
  {L}ojasiewicz inequalities}, Ann. Math. (2015), 221--285.

\bibitem{coldingminicozzisingularset}
\bysame, \emph{The singular set of mean curvature flow with generic
  singularities}, Invent. Math. \textbf{204} (2016), no.~2, 443--471.

\bibitem{jendrejlawrieschlag}
J.~Jendrej, A.~Lawrie, and W.~Schlag, \emph{Continuous in time bubble
  decomposition for the harmonic map heat flow}, Forum of Mathematics, Pi
  \textbf{13} (2025), e4.

\bibitem{leonsimon}
L.~Simon, \emph{Asymptotics for a class of non-linear evolution equations, with
  applications to geometric problems}, Ann. Math. (1983), 525--571.

\bibitem{simonrectifiability}
\bysame, \emph{Rectifiability of the singular set of energy minimizing maps},
  Calculus of Variations and Partial Differential Equations \textbf{3} (1995),
  no.~1, 1--65.

\bibitem{songwaldron}
C.~Song and A.~Waldron, \emph{Harmonic map flow for almost-holomorphic maps},
  Journal of Differential Geometry \textbf{128} (2024), no.~3, 1226--1268.

\bibitem{struwehmfhigherd}
M.~Struwe, \emph{On the evolution of harmonic maps in higher dimensions},
  Journal of differential geometry \textbf{28} (1988), no.~3, 485--502.

\bibitem{toppingwindingbehavior}
P.~M. Topping, \emph{Winding behaviour of finite-time singularities of the
  harmonic map heat flow}, Mathematische Zeitschrift \textbf{247} (2004),
  no.~2, 279--302.

\bibitem{lojasiewiczonS2}
A.~Waldron, \emph{{\L}ojasiewicz inequalities for maps of the 2-sphere},
  ArXiv:2312.16686, submitted.

\bibitem{stricttypeII}
\bysame, \emph{Strict type-{II} blowup in harmonic map flow}, Proceedings of
  the American Mathematical Society \textbf{151} (2023), no.~11, 4893--4907.

\bibitem{zhulojasiewiczcylindrical}
J.~J. Zhu, \emph{{\L}ojasiewicz inequalities, uniqueness and rigidity for
  cylindrical self-shrinkers}, ArXiv:2011.01633.

\bibitem{zhulojasiewiczmeanconvex}
\bysame, \emph{{\L}ojasiewicz inequalities for mean convex self-shrinkers},
  International Mathematics Research Notices \textbf{2023} (2023), no.~2,
  1236--1254.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage{amssymb}
\usepackage{amsmath,amscd}
\usepackage{amsfonts,latexsym,verbatim,amscd,mathrsfs,color,array}
\usepackage[all,cmtip]{xy}
\usepackage{mathrsfs}
\usepackage{multirow}
\usepackage{graphicx}
\usepackage{amsfonts, amsthm}
\usepackage{bbm}
\usepackage[hmargin=1in,vmargin=1in]{geometry}
\usepackage{mathdots}
\usepackage{stmaryrd}
\usepackage{MnSymbol}
\usepackage{bbm}
\usepackage[hidelinks]{hyperref}

\allowdisplaybreaks

\def\XXint#1#2#3{{\setbox0=\hbox{$#1{#2#3}{\int}$ }
		\vcenter{\hbox{$#2#3$ }}\kern-.6\wd0}}

\DeclareMathOperator*{\Osc}{\osc}

\usepackage{accents}
\newlength{\dhatheight}

\newcommand{\osc}{\operatorname{osc}}
\DeclareMathSymbol{\intprod}{\mathbin}{MnSyC}{'270}
\newcommand{\LB}{\left[}
\newcommand{\RB}{\right]}
\newcommand{\LA}{\left\langle}
\newcommand{\RA}{\right\rangle}

\newcommand{\N}{{\mathbb N}}
\newcommand{\C}{{\mathbb C}}

\newcommand{\R}{{\mathbb R}}

\newcommand{\eps}{{\varepsilon}}

\newcommand{\T}{{\mathcal T}}

\newcommand{\p}{{\partial}}

\newcommand{\That}{{ \widehat{\T} }}

\newtheorem{thm}{Theorem}[section]
\newtheorem{lemma}[thm]{Lemma}
\newtheorem*{lemma*}{Lemma}
\newtheorem{prop}[thm]{Proposition}

\newtheorem{cor}[thm]{Corollary}

\newtheorem*{conj*}{Conjecture}

   \newtheoremstyle{others}
     {3pt}
     {2pt}
     {}
     {}
     {\bf}
     {.}
     {.5em}
     {}

\theoremstyle{others}

\newtheorem*{rmk*}{Remark}
\newtheorem{defn}[thm]{Definition}

\numberwithin{equation}{section}

\begin{document}

\title{Weighted \L ojasiewicz inequalities
and regularity of harmonic map flow}
\author{Alex Waldron}
\address{University of Wisconsin, Madison}
\email{waldron@math.wisc.edu}

\begin{abstract}
At a finite-time singularity of harmonic map flow in the critical dimension, we show that a \L ojasiewicz inequality between the quantities appearing in Struwe's monotonicity formula implies continuity of the body map and the no-neck property for bubble-tree decompositions. We prove such an inequality when the target is $S^2,$ yielding both properties in this case.
\end{abstract}

\maketitle

\thispagestyle{empty}



\vspace{-5mm}

\section{Introduction}

\L ojasiewicz(-Simon) inequalities play a key role in the theory of geometric evolution equations; they rule out the notorious non-uniqueness phenomenon for subsequential limits at infinite time \cite{leonsimon}.
In recent years, most prominently in the work of Colding and Minicozzi, \L ojasiewicz inequalities suitable for finite-time singularity analysis have become available.
These have led to uniqueness theorems for the most basic multiplicity-one blowup limits of mean curvature flow, namely, round cylinders \cite{coldingminicozziuniquenesslojasiewicz} and asymptotically conical shrinkers \cite{chodoshschulzeuniquenessofasymptoticallyconical}, as well as mean-convex cylindrical shrinkers \cite{zhulojasiewiczcylindrical, zhulojasiewiczmeanconvex}.

\L ojasiewicz inequalities are also strongly tied to 
questions of regularity. Simon used his classical inequality in the elliptic setting to obtain a strong Hausdorff-dimension estimate and rectifiability of the singular set for area-minimizing currents \cite{simonrectifiability}. Colding and Minicozzi have applied their results to obtain improved regularity of singular sets for mean curvature flow with generic (i.e. cylindrical) singularities at finite time \cite{coldingminicozzisingularset}. They have also demonstrated that differentiability of the ``arrival time'' in mean-convex mean curvature flow depends on a \L ojasiewicz inequality \cite{coldingminicozzidifferentiabilityofthearrivaltime}.

This paper establishes a link between a different regularity question and a \L ojasiewicz inequality. 
The question concerns the so-called ``body map'' $u(T) = \lim_{t \nearrow T} u(t)$ of a solution of harmonic map flow $u : \Sigma \times \LB 0, T \right) \to N$ in the critical dimension, $\dim\left( \Sigma \right) = 2.$ This map is by definition of class $H^1,$ but may fail to be continuous in certain cases, as shown by Topping \cite{toppingwindingbehavior}. Meanwhile, Topping conjectured that if $N$ is real-analytic then the body map must be continuous \cite[p. 286]{toppingwindingbehavior}.

Our main result, Theorem \ref{thm:mainthm} below, is that continuity of the body map follows from the validity of a \L ojasiewicz inequality 
between the Gaussian-weighted quantities appearing in Struwe's monotonicity formula \cite{struwehmfhigherd}. We prove such an inequality, Theorem \ref{thm:firstloj}, in the case that the target manifold is the round 2-sphere, allowing us to 
settle the fundamental case of the above question.\footnote{The continuity question for $N = S^2$ was raised again recently by Jendrej, Lawrie, and Schlag \cite{jendrejlawrieschlag}.} 
Our main estimate also gives the ``no-neck'' property of the bubble-tree decomposition along any sequence approaching the singular time.

The weighted \L ojasiewicz inequality for maps to the 2-sphere is obtained by directly localizing one of the inequalities proved recently by the author \cite{lojasiewiczonS2}.
The localization is enabled by a weighted Poincar\' e inequality, Lemma \ref{lemma:Poincare}, similar to those used in Type-I singularity analysis; this inequality turns out to be particularly strong in the critical dimension. We use the inequality in two places: once in the localization, Theorem \ref{thm:firstloj}, and later in the proof of main theorem, cf. (\ref{psiintegralbound}).

An interesting feature of the proof is that we do not obtain H\"older continuity of the body map, as in the case of a ``strict'' type-II blowup \cite{stricttypeII}, unless the \L ojasiewicz inequality holds with the optimal exponent ($\alpha = 2$). In the case of a fractional exponent, we obtain a $\log$-polynomial oscillation bound which is sufficient for continuity and no necks.

\vspace{5mm}

\section{Weighted \L ojasiewicz inequality}

\subsection{Weighted Poincar{\'e} inequality} Let $(\Sigma, g)$ be a Riemannian surface and $(N,\LA \cdot, \cdot \RA)$ any compact Riemannian manifold. Given a map $u : \Sigma \to N,$ denote its stress-energy tensor by
\begin{equation}\label{stressdef}
S_{ij} = \LA \p_i u, \p_j u \RA - \frac12 g_{ij} |du|_g^2.
\end{equation}
We have
\begin{equation}\label{stressdiv}
\nabla^i S_{ij} = \LA \T (u) , \p_j u \RA,
\end{equation}
see e.g. \cite[(2.5)]{songwaldron}.
Let $X^i$ be any conformal Killing field, satisfying
$$\nabla^{i} X^j + \nabla^j X^i = \mu g^{ij}$$
for some function $\mu.$
Note that $g^{ij} S_{ij} = 0$ in dimension two. Contracting the identity (\ref{stressdiv}) with $X,$ we have
\begin{equation}\label{stressenergyidentity}
\nabla^i \left( X^j S_{ij} \right) = \LA \T (u), X^j \p_j u \RA.
\end{equation}
We now specialize to the case $\Sigma = \R^2$ with the flat metric; the general case requires only small modifications. 
Let $x = x^i \p_i$ be the radial vector field and $r = \sqrt{(x^1)^2 + (x^2)^2}$ the radial coordinate in $\R^2.$ Define the weighted $L^2$-norm
\begin{equation}\label{bartaudef}
\| \alpha \|_\tau = \sqrt{ \int_{\R^2} |\alpha |^2 \, e^{-\frac{r^2}{4\tau} } \, dV }.
\end{equation}
Let
\begin{equation*}\label{Thatpredefn}
\That_\tau (u) := \T(u) - \frac{x}{2 \tau  } \intprod du.
\end{equation*}

\begin{lemma}\label{lemma:Poincare} For any $W^{2,2}$ map $u: \R^2 \to N,$ we have
\begin{equation}\label{twistedtensiondominatesrdu}
\left\| r du \right\|_\tau \leq 4 \tau \left\|\That_\tau(u) \right\|_\tau
\end{equation}
and
\begin{equation}\label{Tdominance}
\| \T(u) \|_{\tau} \leq 3 \left\| \That_\tau(u) \right\|_{\tau}.
\end{equation}
\end{lemma}
\begin{proof}
Integrating (\ref{stressenergyidentity}) by parts against $e^{-\frac{r^2}{4 \tau} },$ 
we obtain
\begin{equation*}
\frac{1}{2\tau} \int x^i x^j S_{ij} e^{-\frac{r^2}{4 \tau} } \, dV = \int \LA \T(u), x^j \p_j u \RA e^{-\frac{r^2}{4 \tau} } \, dV.
\end{equation*}
Unpacking the definition of the stress-energy tensor (\ref{stressdef}), we have
\begin{equation*}
\begin{split}
-  \frac{1}{4 \tau} \int r^2 |du|^2 e^{-\frac{r^2}{4 \tau} } \, dV & = \int \LA \T(u), x^j \p_j u \RA e^{-\frac{r^2}{4 \tau} } \, dV - \int \LA \frac{ x^i  }{ 2\tau } \p_i u, x^j \p_j u \RA e^{-\frac{r^2}{4 \tau} } \, dV \\
& = \int \LA \That_\tau(u), x^j \p_j u \RA e^{-\frac{r^2}{4 \tau} } \, dV.
\end{split}
\end{equation*}
Applying Cauchy-Schwarz, we obtain
$$\frac{1}{4 \tau} \left\| r du \right\|_\tau^2 \leq \left\| \That_\tau(u) \right\|_{\tau} \| x \intprod du \|_\tau.$$
Since $|x \intprod du| \leq r |du|,$ this implies (\ref{twistedtensiondominatesrdu}).
Then (\ref{Tdominance}) follows from (\ref{twistedtensiondominatesrdu}) and the triangle inequality.
\end{proof}

\vspace{5mm}

\subsection{Weighted \L ojasiewicz inequality for maps to $S^2$}

Let
$$\Phi(u) := \frac12 \int |du|^2 e^{ \frac{-r^2}{4} } \, dV$$
and
$$\That(u) : = \That_1(u) = \T(u) - \frac{x}{2 } \intprod du.$$

\begin{thm}\label{thm:firstloj}
Fix $k \in \N$ and $\beta > 0.$ Given any map $u : \R^2 \to S^2$ with $E(u) \leq 4 \pi k,$ there exists $n \in \{0, \ldots, k\}$ such that
$$|\Phi(u) - 4\pi n | \leq C_{k, \beta} \| \That(u) \|_1^{2 - \beta},$$
where $\| \cdot \|_{1}$ is the weighted $L^2$-norm defined by (\ref{bartaudef}).
\end{thm}
\begin{proof} 

By the weighted Poincar\'e inequality of Lemma \ref{lemma:Poincare}, we have
$$ \int r^2 |du|^2 e^{-\frac{r^2}{4}}  + \| \T(u) \|^2_{L^2(B_2)} \leq C \| \That(u) \|_1^2.$$
In particular, this gives $ E(u, B_2 \setminus B_1) \leq C\| \That(u) \|_1^2.$ We can assume without loss of generality that $ \| \That(u) \|^2_1 \leq \eps_0 / C.$ 

Let $0 \leq \lambda \leq 1$ be the smallest radius such that
$$\int_{B_2 \setminus B_\lambda} |du|^2 \leq \eps_0.$$
We have
$$\lambda^2 \int_{B_2 \setminus B_\lambda} |du|^2 \leq \int r^2 |du|^2 e^{-\frac{r^2}{4}} \, dV \leq  C \| \That(u) \|_1^2.$$
This gives
$$\lambda \leq  C \| \That(u) \|_1.$$
Now, \cite[Corollary 1.3]{lojasiewiczonS2} reads
\begin{equation}\label{quantitativeloj:cestimate}
\begin{split}
&  \left| E(u, B_1) - 4 \pi n \right| \\
 & \qquad  \leq C_{k,\beta} \left( \| \T(u) \|_{L^2(B_2)}^2 + E(u, B_{2} \setminus B_1) + \lambda^{1 - \beta} \left( \| \T(u) \|_{L^2(B_2)} + \sqrt{E(u, B_{2} \setminus B_1)} \right) \right).
 \end{split}
\end{equation}
Inserting the above estimates into the RHS of (\ref{quantitativeloj:cestimate}), we get
$$\left| E(u, B_1) - 4 \pi n \right| \leq C \| \That(u) \|_1^{2 - \beta}.$$
We further have
$$\left| E(u, B_1) - \frac12 \int_{B_1} |du|^2 e^{\frac{-r^2}{4}} \right| \leq \int_{B_1} r^2 |du|^2 \leq C \| \That(u) \|_1^{2}.$$
Finally, we have
$$\int_{\R^2 \setminus B_1} |du|^2 e^{\frac{-r^2}{4}} \leq \int_{\R^2 \setminus B_1} r^2 |du|^2 e^{\frac{-r^2}{4}} \leq C \| \That(u) \|_1^{2}.$$
Combining these, we obtain the desired estimate.
\end{proof}

\vspace{5mm}

\section{Monotonicity formulae in dimension two}

Recall the \emph{pointwise energy identity} for harmonic map flow,
\begin{equation}\label{pointwiseenergy}
\frac12 \frac{\p }{\p t} |du|^2 + |\T(u)|^2 = \nabla^i \nabla^j S_{ij},
\end{equation}
see \cite[(2.7)]{songwaldron}.

\begin{lemma} 
Let
$$w = w(r,t) = \exp \left( -{\frac{\phi(x)}{2\tau(t)}} \right).$$
For a solution $u : \Sigma \to N$ of harmonic map flow, we have
\begin{equation*}
\begin{split}
\frac12 \frac{\p}{\p t} \left( |du|^2 w \right) + \left| \T(u) - \frac{\nabla \phi}{2 \tau} \intprod du \right|^2 w & = \nabla^i \nabla^j \left( S_{ij} w \right) \\
& \qquad + \frac{w}{2\tau} \nabla^i \nabla^j \phi S_{ij} + \frac{\left( |\nabla \phi|^2 + 2 \phi \tau'(t) \right) w}{8\tau^2}  |du|^2.
\end{split}
\end{equation*}
\end{lemma}
\begin{proof}
Multiplying the pointwise energy identity (\ref{pointwiseenergy}) by $w,$ 
we obtain
\begin{equation*}
\begin{split}
\frac12 \frac{\p}{\p t} \left( |du|^2 w \right) - \frac{\tau'(t) }{4\tau^2} \phi |du|^2 w + |\T(u)|^2 w & = \nabla^i \nabla^j \left( S_{ij} w \right) - 2 \nabla^i w \nabla^j S_{ij} - \nabla^i \nabla^j w S_{ij} \\
& = \nabla^i \nabla^j \left( S_{ij} w \right) + 2\LA \T(u) , \frac{\nabla \phi \intprod du}{2\tau}  \RA w \\
&  \quad - \frac{ \left(2\tau \nabla^i \nabla^j \phi + \nabla^i \phi \nabla^j \phi \right) }{4\tau^2} S_{ij} w.
\end{split}
\end{equation*}
Recalling the definition of the stress-energy tensor (\ref{stressdef}), we can rearrange to obtain the desired formula.
\end{proof}

We have the following versions of Struwe's monotonicity formula in the critical dimension.

\begin{thm}\label{thm:monotonicities} Let $\Sigma = \R^2$ and take
$$\Phi_{\tau}(u) = \frac12 \int |du|^2 e^{-\frac{r^2}{4 \tau}} \, dV, \quad \Psi_{\tau}(u) = \frac12 \int r^2 |du|^2 e^{-\frac{r^2}{4 \tau}} \, dV.$$
We have
\begin{equation}\label{Phimonotonicity}
\frac{d}{dt} \Phi_{\tau}(u(t)) + \| \That_{\tau}(u(t)) \|_{\tau}^2 = \frac{1 + \tau'(t)}{4 \tau^2} \Psi_\tau(u(t))
\end{equation}
and
\begin{equation}\label{Psimonotonicity}
\frac{d}{dt} \Psi_{\tau}(u(t)) + \| r \That_{\tau}(u(t)) \|_{\tau}^2 = \frac{1 + \tau'(t)}{4 \tau^2} \int r^4 |du|^2 e^{-\frac{r^2}{4 \tau}} \, dV.
\end{equation}
\end{thm}
\begin{proof} Take $\phi = \frac{r^2}{2}$ in the previous Lemma, so that $w = e^{\frac{-r^2}{4\tau}}.$
We have
\begin{equation}\label{pointwisemonotR2}
\begin{split}
\frac12 \frac{\p}{\p t} \left( |du|^2 w \right) + \left| \That_{\tau}(u) \right|^2 w & = \nabla^i \nabla^j \left( S_{ij} w \right) + \frac{\left( 1 + \tau'(t) \right)r^2}{4 \tau^2} |du|^2 w.
\end{split}
\end{equation}
The identity (\ref{Phimonotonicity}) follows by integrating (\ref{pointwisemonotR2}) in space. The identity (\ref{Psimonotonicity}) follows by integrating by parts against $r^2$ and using $\nabla^i\nabla^j r^2 = 2 \delta^{ij}$ together with tracelessness of $S_{ij}.$
\end{proof}

\vspace{5mm}

\section{Estimates on energy density and oscillation}

\begin{lemma}\label{lemma:rduestimate}
Fix $0 < 2 \lambda \leq R \leq \sqrt{T}$ and $\kappa > 0.$ 
Let $u : \R^2 \times \LB 0, T \right) \to N$ be a classical solution of harmonic map flow.
Suppose that $R^2 \leq T_1 \leq T$ is such that 
\begin{equation*}
\sup_{T_1 - R^2 \leq t \leq T_1 - \lambda^2} \frac{\Psi_{T_1 - t} \left( u(t) \right)}{T_1 - t} \leq \eps_0,
\end{equation*}
where $\eps_0 > 0$ depends on the geometry of $N.$ Letting
\begin{equation*}
\eta : = \sqrt{ \frac{\Psi_{4R^2} \left( u(T_1 - R^2) \right)}{R^2} } 
\end{equation*}
and
 \begin{equation}\label{psidefn}
\psi \left( \log \frac{R}{\sqrt{T_1 - t} } \right) : = \sqrt{\frac{\Psi_{T_1 - t} \left( u(t) \right)}{T_1 - t} },
\end{equation}
we have
\begin{equation*}
r |du(x,t)| \leq C_{\kappa} \left( \int_0^{\log \frac{R}{r}} \psi\left( \sigma \right) e^{\kappa \left( \sigma - \log \frac{R}{r} \right)} \, d\sigma  + \eta \left( \frac{r}{R} \right)^\kappa \right)
\end{equation*}
for 
$\lambda \leq r \leq \min \left\{ \left( C_{N, \kappa} \sqrt[\kappa]{\eta} \right)^{-1} , 1 \right\} R$ and
$T_1 - \frac14 r^2 \leq t < T_1.$ 
\end{lemma}
\begin{proof}
We may assume without loss of generality that $R = T_1 = 1.$
Let
$$\bar{\psi}(t) = \sqrt{ \frac{\Psi_{4(1 - t)}(u(t))}{1 - t} }.$$
We claim that
\begin{equation}\label{psiclaim}
\bar{\psi}(t) \leq \eta (1 - t)^{\frac{\kappa}{2} } + \frac{e^{4 + 4\kappa} }{\kappa} \int_0^{- \frac12 \log \left( 1 - t \right) } \psi \left( s \right)e^{\kappa \left(s + \frac12 \log \left( 1 - t \right) \right)} \, ds
\end{equation}
for all $0 \leq t < 1.$

Applying the previous theorem with $\tau(t) = 4(1 - t),$ we obtain
\begin{equation}\label{psi'(t)}
\begin{split}
\frac{d \bar{\psi}^2(t) }{dt} 
& \leq \frac{1}{1 - t} \left( \bar{\psi}^2(t) - \frac{3}{4 \tau^2} \int r^4 |du|^2 e^{-\frac{r^2}{4\tau}} \, dV \right)
\end{split}
\end{equation}
Note that since $e^{a-3x} > a - 3x,$ we have 
$$3x e^{-x} > a e^{-x} - e^{a - 4x}.$$
Taking $a = 4 + 4\kappa$ and $x = \frac{r^2}{4 \tau},$ and multiplying by $-r^2,$ we obtain
$$-\frac{3r^4}{4 \tau} e^{-\frac{r^2}{4 \tau} } < r^2 \left( e^{4 + 4 \kappa -\frac{r^2}{\tau} } - (4 + 4 \kappa) e^{-\frac{r^2}{4 \tau} } \right).$$
Using this in the integrand of the second term on the RHS of (\ref{psi'(t)}), we get
\begin{equation*}
\begin{split}
(1 -t) \frac{d\bar{\psi}^2(t)}{dt} & \leq \bar{\psi}^2(t) + e^{4 + 4 \kappa} \frac{\Psi_{1 - t}(u(t))}{\tau} - (1 + \kappa)\bar{\psi}^2(t) \\
& \leq e^{4 + 4\kappa} \psi^2 \left( - \frac12 \log \left( 1 - t \right) \right) - \kappa \bar{\psi}^2(t).
\end{split}
\end{equation*}
Let $s = - \frac12 \log \left( 1 - t \right),$ so that $\frac{d}{ds} = 2 (1 - t) \frac{d}{dt}$ and the above reduces to
$$\frac{d \bar{\psi}^2(s)}{ds} \leq 2\left( e^{4 + 4\kappa} \psi^2(s) - \kappa \bar{\psi}^2(s) \right).$$
Since $\psi(s) \leq \bar{\psi}(s),$ we may use the product rule and divide out by $\bar{\psi}(s),$ to obtain
$$\frac{d \bar{\psi}(s)}{ds} \leq e^{4 + 4\kappa} \psi(s) - \kappa \bar{\psi}(s) .$$
By Gronwall's inequality, we have
\begin{equation*}
\begin{split}
\bar{\psi}(s) & \leq \left( \bar{\psi}(0)  + e^{4 + 4\kappa} \int_0^s \psi(\sigma) e^{ \kappa \sigma} \, d\sigma \right) e^{-\kappa s} \\
& \leq \eta (1 - t)^{\kappa/2} + e^{4 + 4\kappa} \int_0^s \psi(s) e^{\kappa (\sigma - s) } \, d \sigma. 
\end{split}
\end{equation*}
This establishes the claim (\ref{psiclaim}).

Now, let $\rho \in \LB \lambda, R \RB$ and take $\tilde{\tau}(t) = 3\rho^2 + 1 - t,$ so that $\tilde{\tau}(1 - \rho^2) = 4 \rho^2$ and $\tilde{\tau}'(t) = -1.$ 
Integrating (\ref{Psimonotonicity}) from $1 - \rho^2$ to $t \in \LB 1 - \rho^2, 1 \right),$ we obtain
$$\Psi_{\tilde{\tau}(t)}(u(t)) \leq \Psi_{\tilde{\tau}(1 - \rho^2)} \left( u \left( 1 - \rho^2 \right) \right) = \rho^2 \bar{\psi}^2(1 - \rho^2). $$ 
Since $3 \rho^2 \leq \tilde{\tau}(t) \leq 4 \rho^2,$ we get
\begin{equation}\label{dusquaredpreestimate}
\begin{split}
\int_{\frac12\rho }^{2 \rho } |du(t)|^2\, dV & \leq C \int_{\frac12 \rho }^{2 \rho } \frac{r^2}{\tilde{\tau}(t)} |du(t)|^2 e^{- \frac{r^2}{ 4 \tilde{\tau}(t) } } \, dV  \leq C \frac{\Psi_{\tilde{\tau}(t)}(u(t))}{\rho^2} \leq \bar{\psi}^2(1 - \rho^2).
\end{split}
\end{equation}
Assuming that $\rho^{2\kappa} \leq \frac{\eps_0}{\eta^2},$ (\ref{psiclaim}) implies that the RHS of (\ref{dusquaredpreestimate}) is less than $C \eps_0.$ We may apply standard epsilon-regularity, see e.g. \cite[Theorem 3.4]{songwaldron}, and (\ref{psiclaim}), to obtain the desired estimate.
\end{proof}

\begin{thm}\label{thm:mainthm}
Let $u : \R^2 \times \LB 0, T \right) \to N$ be a classical solution of harmonic map flow. Suppose that $0 < T_1 \leq T$ and $0 < R \leq \sqrt{T_1}$ are such that 
\begin{equation*}
\Psi_{4R^2} \left( u(T_1 - R^2) \right) \leq \eta R^2.
\end{equation*}
Suppose that there exists $E_0 \geq 0,$ $0 < \eps \leq \eps_0,$ and $0 < 2\lambda \leq R$ such that
\begin{equation}\label{mainthm:Phiassn}
\Phi_{R^2}(u(T_1 - R^2)) \leq E_0 + \eps, \qquad \Phi_{\lambda^2}(u(T_1 - \lambda^2)) \geq E_0 - \eps,
\end{equation}
and,
for all $T_1 - R^2 \leq t < T_1 - \lambda^2,$ $u(t)$ obeys a \L ojasiewicz inequality of the form
\begin{equation}\label{mainthm:lojasiewiczassumption}
| \Phi_{T_1 - t}(u(t)) - E_0| \leq \left( K \sqrt{T_1 - t} \| \That_{T_1 - t}(u(t)) \|_{T_1 - t} \right)^\alpha
\end{equation}
for some $K \geq 1$ and $1 < \alpha \leq 2.$ 
Then there exists a constant $C_{\ref{thm:mainthm}a},$ depending on $N, K,$ and $\alpha,$ as well as $C_{\ref{thm:mainthm}b},$ depending on $N$ and $K,$ as follows.

\vspace{2mm}

\noindent (a) If $\alpha < 2,$ we have
\begin{equation*}
\Osc_{ \frac{\lambda R}{r} \leq |x| \leq r} u(x,t) \leq C_{\ref{thm:mainthm}a} \left( \left( \eps^{\frac{\alpha - 2}{\alpha} } + \log \frac{R}{r} \right)^{\frac{\alpha - 1}{\alpha - 2} } + \eta \frac{ r}{R} \right) 
\end{equation*}
for $\sqrt{ R \lambda } \leq r \leq \min \left\{ 1, (C_{\ref{thm:mainthm}a} \eta)^{-1} \right\}R $ and $T_1 - \frac14 r^2 \leq t < T_1.$

\vspace{2mm}

\noindent (b) If $\alpha = 2,$ we have
\begin{equation*}
\Osc_{ \frac{\lambda R}{r} \leq |x| \leq r}  u(x,t) \leq C_{\ref{thm:mainthm}b} \left( \eps \left( \frac{r}{R} \right)^{\frac{1}{K}} + \eta \frac{ r}{R} \right) 
\end{equation*}
for $\sqrt{ R \lambda } \leq r \leq \min \left\{ 1, (C_{\ref{thm:mainthm}b} \eta)^{-1} \right\}R $ and $T_1 - \frac14 r^2 \leq t < T_1.$
\end{thm}
\begin{proof}
Let
$$\phi(t) := \Phi_{T_1 - t}(u(t)) - E_0.$$
By (\ref{Phimonotonicity}) and the \L ojasiewicz inequality (\ref{mainthm:lojasiewiczassumption}), we have
\begin{equation*}
\frac{d \phi(t)}{dt} + \frac{|\phi(t)|^{\frac{2}{\alpha} } }{K^2 (T_1 - t) } \leq 0.
\end{equation*}
Letting $s = -\frac12 \log \frac{T_1 - t}{R^2}$ as in the last proof, we get
\begin{equation}
\frac{d \phi(s)}{ds} + \frac{2 }{K^{2}} |\phi(s)|^{\frac{2}{\alpha} } \leq 0.
\end{equation}
Let
$$\varphi(s) : = \left( \eps^{-\frac{2 - \alpha}{\alpha} } + \frac{2(2 - \alpha)}{K^{2} \alpha} s \right)^{-\frac{\alpha}{2 - \alpha} },$$
so that $\varphi(0) = \eps$ and
$$\frac{d\varphi(s)}{ds} = - \frac{2}{K^{2}} \varphi(s)^{\frac{2}{\alpha}}$$
for $s \geq 0.$ By Gronwall's inequality, we have
\begin{equation}\label{phiboundfromabove}
\phi(s) \leq \varphi (s).
\end{equation}
Meanwhile, we also have
$$\frac{d\left( - \varphi \left( \log \frac{R}{\lambda} - s \right) \right)}{ds} = - \frac{2}{K^{2}} \varphi \left( \log \frac{R}{\lambda} - s \right)^{\frac{2}{\alpha}}$$
for $s \leq \log \frac{R}{\lambda}.$
Applying Gronwall's inequality backwards from $s = \log \frac{R}{\lambda}$ to $s = 0,$ we obtain
\begin{equation}\label{phivarphibound}
-\varphi \left( \log \frac{R}{\lambda} - s \right) \leq \phi(s) \leq \varphi (s).
\end{equation}
Next, we translate this into an integral bound on $\Psi.$ Let
$$\delta(s) := \sqrt{T_1 - t} \| \That_{T_1 - t} \|_{T_1 - t},$$
where $T_1 - t = R^2 e^{-2s}.$
By (\ref{Phimonotonicity}), we have
$$\frac{d}{ds} \phi(s)^{\frac{\alpha - 1}{\alpha}} = - \frac{\alpha - 1}{\alpha} |\phi(s)|^{-\frac{1}{\alpha}} \delta^2(s),$$
where we define $\phi(s)^{\frac{\alpha - 1}{\alpha}} = -|\phi(s)|^{\frac{\alpha - 1}{\alpha}}$ if $\phi(s)$ is negative.
Let $S = \log R / \lambda.$ For $s \leq S/2,$ applying the \L ojasiewicz inequality (\ref{mainthm:lojasiewiczassumption}), we may write
\begin{equation*}
\begin{split}
\int_{s}^{S - s} \delta(\sigma) \, d\sigma = \int_{s}^{S - s} \delta^2(\sigma) \delta(\sigma)^{-1} \, d\sigma & \leq \int_{s}^{S - s} \delta(\sigma)^2 |\phi(\sigma)|^{-\frac{1}{\alpha}} \, d\sigma \\
& \leq - \frac{\alpha}{\alpha - 1} \int_{s}^{S - s} \frac{d}{d\sigma} \phi(\sigma)^{\frac{\alpha - 1}{\alpha}} \, d\sigma \\
& \leq \frac{\alpha}{\alpha -1} \left( \phi(s)^{\frac{\alpha -1}{\alpha}} - \phi(S - s)^{\frac{\alpha -1}{\alpha}} \right) \\
& \leq \frac{2\alpha}{\alpha -1} \varphi(s)^{\frac{\alpha -1}{\alpha}},
\end{split}
\end{equation*}
where we have used (\ref{phivarphibound}).

Define $\psi(s)$ by (\ref{psidefn}). Applying Lemma \ref{lemma:Poincare}, we obtain
\begin{equation}\label{psiintegralbound}
\begin{split}
\int_{s}^{S - s} \psi(\sigma) \, d\sigma \leq 4 \int_{s}^{S - s} \delta(\sigma) \, d\sigma \leq \frac{8\alpha}{\alpha -1} \varphi(s)^{\frac{\alpha -1}{\alpha}}.
\end{split}
\end{equation}
We now use (\ref{psiintegralbound}) to estimate the oscillation of $u.$ Let $s = \log R / r.$ We have
$$\Osc_{ \frac{\lambda R}{r} \leq |x| \leq r}  u(x,t) \leq \int^r_{\lambda R / r} |du| \, d\rho = \int_s^{S - s} r |du| \, d\varsigma.$$
Applying Lemma \ref{lemma:rduestimate} with $\kappa = 1,$ for $T_1 - \frac14 r^2 \leq t < T_1,$ we get
\begin{equation}\label{oscUestimate}
\begin{split}
\Osc_{ \frac{\lambda R}{r} \leq |x| \leq r}  u(x,t) \leq \int_s^{S - s} r |du| \, d\varsigma & \leq C \int_s^{S - s} \left( \int_0^{\varsigma} \psi\left( \sigma \right) e^{ \sigma - \varsigma } \, d\sigma  + \eta e^{-\varsigma} \right) \, d\varsigma \\
& \leq C \left( \int_s^{S - s} \!\!\! \int_0^{\varsigma} \psi\left( \sigma \right) e^{ \sigma - \varsigma } \, d\sigma \, d\varsigma  + \eta e^{-s} \right).
\end{split}
\end{equation}
Changing variables $\nu = \varsigma - \sigma,$ we can estimate the first integral by
\begin{equation*}
\begin{split}
\int_s^{S - s} \!\!\! \int_0^{\varsigma} \psi\left( \sigma \right) e^{\sigma - \varsigma} \, d\sigma \, d\varsigma & = \int_{0}^{S - s} e^{-\nu} \int_{\max \{s - \nu, 0\} }^{S - s - \nu} \psi(\sigma ) \, d\sigma d\nu \\
& \leq C \int_{0}^{S - s} e^{- \nu} \varphi \left( \max \{s - \nu, 0 \} \right)^{\frac{\alpha - 1}{\alpha}} d\nu \\
& \leq C \varphi(s )^{\frac{\alpha - 1}{\alpha}},
\end{split}
\end{equation*}
where we have used (\ref{psiintegralbound}). Returning to (\ref{oscUestimate}), we obtain the bound of (a). The bound (b) is proved similarly.
\end{proof}

\begin{cor}\label{cor:oscillation}
Take $T_1 = T$ in the previous theorem and suppose that the assumptions (\ref{mainthm:Phiassn}-\ref{mainthm:lojasiewiczassumption}) hold for all $\lambda \in \left( 0, R \RB.$ Then the limit $u(x,T) : = \lim_{t \nearrow T} u(x,t)$ exists for all $x \in B_{R/2} \setminus \{0\},$ with
\begin{equation*}
\Osc_{B_r} u(T) = O \left| \log r \right|^{\frac{\alpha - 1}{\alpha - 2}}
\end{equation*}
as $r \searrow 0$ if $1 < \alpha < 2,$ and is H\"older continuous if $\alpha =2.$
\end{cor}
\begin{proof} This follows by letting $\lambda \searrow 0$ in the previous theorem.
\end{proof}

\begin{cor} 
Let $u: \R^2 \times \LB 0, T \right) \to S^2$ be a classical solution of harmonic map flow. 

\vspace{2mm}

\noindent (a) The body map $u(T)$ extends to a continuous map $\R^2 \to S^2.$

\vspace{2mm}

\noindent (b) Given a point $p \in \R^2$ and any sequence of times $t_i \nearrow T,$ the maps $u(t_i)$ sub-converge in the bubble-tree sense without necks. Specifically, there exist finitely many nonconstant harmonic maps $\phi_k : \R^2 \to S^2,$ obtained as rescaled limits of $u(t_i),$ such that the energy identity holds:
$$\lim_{r \searrow 0} \lim_{t \nearrow T} E(u(t), B_r(p) ) = \sum_k E(\phi_k).$$
Moreover, we have $\lim_{x \to \infty} \phi_k(x) = u(p,T)$ for at least one $k.$

\end{cor}
\begin{proof} ($a$) It suffices to prove continuity of $u(T)$ at $0 \in \R^2.$ By monotonicity of $\Phi,$ the limit
$$E_0 : = \lim_{R \searrow 0} \Phi_{R^2} \left( u(T - R^2) \right) \geq 0$$
exists, and we may take $R > 0$ such that
$$\Phi_{R^2}(u(T - R^2)) \leq E_0 + \eps_0.$$
Now, the \L ojasiewicz inequality of Theorem \ref{thm:firstloj} gives (\ref{mainthm:lojasiewiczassumption}) after rescaling by $\tau = T - t,$ where we must have $E_0 = 4 \pi n.$ 
The result follows from the previous Corollary.

\vspace{2mm}

\noindent ($b$) The proof is standard in light of Corollary \ref{cor:oscillation} (see \cite[Theorem 6.8]{songwaldron} for a recent account), and we omit it.
\end{proof}

\bibliographystyle{amsinitial}
\bibliography{biblio}

\end{document}